OpenStudy (anonymous):

if f(x)=(x^2-8x+9)(x^2-6x+13) has four complex solutions written in factored form as f(x)=(a+bi)(a-bi)(c+di)(c-di). What is the sum of a, b, c, and d?

OpenStudy (anonymous):

no, it should be \[ f(x)=(x-(a+bi))(x-(a-bi))(x-(c+di))(x-(c-di))\]

OpenStudy (anonymous):

you can solve both of these equations easily by completing the square first one is \[x^2-8x+9=0\] \[x^2-8x=-9\] \[(x-4)^2=-9+16=7\] \[x-4=\pm\sqrt{7}\] \[x=4\pm\sqrt{7}\] both roots are real not complex

OpenStudy (anonymous):

this one \[x^2-6x+13=0\] has complex roots, they are \[x^2-6x=-13\] \[(x-3)^2=-13+9=-4\] \[x-3=\pm 2i\] \[x=3\pm2i\]

OpenStudy (anonymous):

you can still add up the roots, you get \(3+3+4+4=14\)

OpenStudy (anonymous):

Thank You :)